OpenStudy (anonymous):

Prove the statement using the the epsilon and delta definition of limit. lim x/5= 3/5 x->3

OpenStudy (amistre64):

f(x)-L < e ; |x-3|<d

OpenStudy (anonymous):

lim x-> 3 means for all delta > 0, there exists an epsilon > 0 such that |x-3| < epsilon then |x/5-3/5| < delta from |x/5-3/5| we see that (-1)delta < (x-3)/5 < delta multiply through by 5 to get |(x-3)| < (5)x(delta) and as stated earlier |x-3| < epsilon so epsilon < 5delta

OpenStudy (amistre64):

x/5 = 3/5 ; *5 x = 3 x-3

OpenStudy (anonymous):

Why is it -3?

OpenStudy (amistre64):

its just the definition as x->c

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

when x-c = 0 that means you reach the spot

OpenStudy (anonymous):

from the definition |x-a| there there is |f(x)-L| < delta

OpenStudy (amistre64):

since we dont care about the value at x=c; as the distance between x and c gets closer and closer; we get a delta of |x-c|

OpenStudy (anonymous):

ah ok, too many freakin laws to remember

OpenStudy (amistre64):

:) yeah, but if you can see them for what they are, they are simpler. spose your at x=2 and wherever I am is c=10 as i get closer to you, the distance between us is measured by |x-c|; |2-10| = 8 i sneak closer to you and im at c=4, the distance between us is: |x-c| = |2-4| = 2; and so the measure of delta can be between 0 and x=c or there about

OpenStudy (amistre64):

err... \[0<|x-c|<\delta\]